OpenStudy (anonymous):

write an equation of the line that passes through the given points. (2,-4),(0,-4)

OpenStudy (anonymous):

Find the slope Slope=y2−y1x2−x1 and then use the slope and one of the given points and use the point-slope formula to get the equation of the line. y−y1=m(x−x1)

OpenStudy (anonymous):

IS THAT THE ANSWER? I WANT THE ANSWER?

OpenStudy (anonymous):

It's an explanation of how you can find the answer. What's the first step, then?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

Start out computing the slope. Slope=(y2−y1)/(x2−x1)

OpenStudy (anonymous):

okae NEXT?

OpenStudy (anonymous):

The slope is called 'm'. then use the slope and one of the given points and use the point-slope formula to get the equation of the line. y−y1=m(x−x1)

OpenStudy (anonymous):

ok i kinda got it now :)

OpenStudy (anonymous):

great!

OpenStudy (anonymous):

ikr haha